### v1

@@ 4.2 Comparison with PFMs of similar parameter scale
We compare the proposed StainNet with five PFMs of comparable parameter size: CTransPath [wang2022transformer], PathoDuet [hua2024pathoduet], HIPT [chen2022scaling], Lunit [kang2023benchmarking], and ResNet-50 (ImageNet pre-trained) [he2016deep]. For the WSI task, we evaluate six MIL models: (1) ABMIL [ilse2018attention], a gated-attention MIL,(2) SiMLP [li2025can], a method combining unsupervised average pooling with an MLP classifier, (3) TransMIL [shao2021transmil], a transformer-based MIL, (4) WiKG [li2024dynamic], a dynamic graph representation method, (5) AMDMIL [ling2024agent], a method combining masked denosing with agent tokens, and (6) S4MIL [fillioux2023structured], a state space model-based MIL. For the ROI tasks, we freeze each PFM and apply two probe settings: a linear probe and an MLP probe for fine-tuning. Table 1 and Table 2 respectively show the results on the slide-level NTUH-Ki67-Liver task and ROI-level BCI and MIST tasks.
For NTUH-Ki67-Liver, StainNet consistently achieves the best overall performance across all MIL encoders. We also observe that lightweight MIL encoders exacerbate the performance gap between H&E-stained image pre-trained PFMs, whereas more complex MIL encoders mitigate this gap. For example, StainNet outperforms the second-best PFM in terms of balanced accuracy by 5.4% with ABMIL, whereas the margin narrows to 2.5% when using WiKG. For BCI and MIST, StainNet continues to outperform significantly, especially on MIST, where StainNet surpasses the popular CTransPath by 11.7% and 15.4% in balanced accuracy under the linear and MLP probe settings.
@@
| BCI 4 classes datasets | Acc. | Bal acc. | AUC | F1-score | Acc. | Bal acc. | AUC | F1-score |
| ResNet-50 | ||||||||
| CTransPath | ||||||||
| PathoDuet | ||||||||
| HIPT | ||||||||
| Lunit | ||||||||
| StainNet (Ours) | ||||||||
@@
| MIST 4 classes datasets | Acc. | Bal acc. | AUC | F1-score | Acc. | Bal acc. | AUC | F1-score |
| ResNet-50 | ||||||||
| CTransPath | ||||||||
| PathoDuet | ||||||||
| HIPT | ||||||||
| Lunit | ||||||||
| StainNet (Ours) | ||||||||
4.3 Comparison with larger PFMs
We further compare StainNet with three recently proposed larger PFMs: PathOrchestra [yan2025pathorchestra], GPFM [ma2025generalizable], and UNI [chen2024towards] as shown in Figure 3. Although StainNet has nearly 14 fewer parameters than them (22M vs 303M), it still achieves the best performance. For example, StainNet surpasses these three large PFMs by 3.72%, 4.70%, and 11.15% in terms of balanced accuracy (80.09% vs. 76.37%, 75.39%, and 68.94%). These results demonstrate that representations learned from large-scale special-stain images provide substantial advantages over H&E-based PFMs when applied to special-stain tasks.
@@ 4.4 Comparison of image retrieval
We also examine the retrieval capability of StainNet to assess its semantic representation quality on special staining histology images. We compare StainNet with CTransPath and UNI on the MIST dataset, as shown in Figure 4. We observe that StainNet consistently achieves the best Recall@K and mAP@K across all choices of K, indicating that it has learned a well-structured embedding space for special staining images.
@@ 4.5 Effectiveness of different fine-tuning data ratios
We further evaluate the fine-tuning performance of StainNet under varying amounts of downstream special-stain training data to assess its learning capability. Figure 5 presents the results of StainNet, CTransPath, and UNI on the MIST dataset using linear and MLP probes under different proportions of fine-tuning samples. We observe that although StainNet performs slightly worse than UNI under the 10% data setting with a linear probe, it achieves the best performance in all other settings and substantially outperforms the CTransPath of similar parameter size. Notably, applying an MLP probe significantly increases the performance gap between StainNet and UNI in low-fine-tuning data scenarios.
4.6 Performance variation under different pre-training iterations
Finally, we investigate the performance variation of StainNet under different pre-training iterations and the behavior of its loss curve. We believe these observations can provide valuable insights for the development of SSL methods in computational pathology, where pre-training details such as the expected loss curve and convergence behavior are often underreported. Figure 6 illustrates the pre-training loss curve over 100 epochs and the corresponding linear probe performance on the MIST dataset at six iteration checkpoints. We observe that the loss plateaus between epoch 30 and 50, but with significant performance improvement during this period. After epoch 60, the overall loss begins to decrease again, but with slight changes in performance. Ultimately, the model reaches its optimal state at the final iteration. These findings suggest that the shift in the DINO pre-training loss may serve as a valuable indicator of representation quality. While additional exploration is required to generalize these observations to other settings, we encourage future SSL development in computational pathology to consider allocating more training epochs, which may be crucial for achieving optimal model performance.
@@ 5 Conclusion and future works
In this work, we propose StainNet, a lightweight pathology foundation model specifically for special staining histopathology image representation. Utilizing the DINO self-supervised framework, StainNet is trained from over 1.4 million unlabeled images. We evaluate StainNet on one in-house WSI task and two public ROI-level tasks, including classification, few-ratio settings, and image retrieval. The results demonstrate that StainNet learns strong and robust representations for special staining pathology images.
For future work, we plan to develop a larger StainNet by expanding the pre-training dataset, scaling the model parameter size, and evaluating it on broader downstream tasks. We will also explore integrating StainNet with H&E-based PFMs for multimodal applications such as survival prognosis and therapy response prediction. We believe that StainNet will further accelerate the research and applications in computational pathology involving diverse histological modalities.

### v2

@@ 4.2 Comparison results on WSI tasks
For the WSI task, we compare the proposed StainNet-Small and StainNet-Base with 12 PFMs from small to large parameter size: (1) ResNet-50 (ImageNet pre-trained) [he2016deep], (2) CTransPath [wang2022transformer], (3) PathoDuet [hua2024pathoduet], (4) HIPT [chen2022scaling], (5) Lunit [kang2023benchmarking], (6) UNI [chen2024towards], (7) PathOrchestra [yan2025pathorchestra], (8) GPFM [ma2025generalizable], (9) UNI-2 [chen2024towards], (10) Virchow-2 [zimmermann2024virchow2], (11) Prov-GigaPath [xu2024whole], and (12) H-optimus-0 [hoptimus0]. For the WSI task, we evaluate six MIL models: (1) ABMIL [ilse2018attention], a gated-attention MIL, (2) SiMLP [li2025can], a method combining unsupervised average pooling with an MLP classifier, (3) TransMIL [shao2021transmil], a transformer-based MIL, (4) WiKG [li2024dynamic], a dynamic graph representation method, (5) AMDMIL [ling2024agent], a method combining masked denosing with agent tokens, and (6) S4MIL [fillioux2023structured], a state space model-based MIL. Table 1 shows the comparison results of the StainNet series models on the six MIL models on NTUH-Ki67-Liver.
Table 2 shows the comparison results of ABMIL on P53-UCEC and MLH1-UCEC. We observe that the StainNet-Base achieves overall best performance on the three WSI tasks. We also report the paired t-test p-values between StainNet-Base and other PFMs on the P53-UCEC dataset, which are shown in Appendix C. For example, when compared with billion-scale models such as Prov-GigaPath and H-optimus-0, the 86M-parameter StainNet-Base still outperforms them on NTUH-Ki67-Liver by 0.83% and 3.38%. Moreover, it outperforms them by 10.9% and 15.2% in balanced accuracy on MLH1-UCEC. It is noteworthy that although StainNet-Small has extremely small parameters (only 22M), it remains highly competitive. For instance, on NTUH-Ki67-Liver, it achieves overall best performance compared to PFMs with similar parameter size, and outperforms UNI and GPFM (303M) by 0.13% and 1.38%. These results demonstrate the advantages of large-scale pre-training on non-H&E pathology images.
@@
| BCI | Acc. | Bal acc. | AUC | F1-score | Acc. | Bal acc. | AUC | F1-score |
| ResNet-50 [he2016deep] | ||||||||
| CTransPath [wang2022transformer] | ||||||||
| PathoDuet [hua2024pathoduet] | ||||||||
| HIPT [chen2022scaling] | ||||||||
| Lunit [kang2023benchmarking] | ||||||||
| UNI [chen2024towards] | ||||||||
| PathOrchestra [yan2025pathorchestra] | ||||||||
| GPFM [ma2025generalizable] | ||||||||
| UNI-2 [chen2024towards] | ||||||||
| Virchow-2 [zimmermann2024virchow2] | ||||||||
| Prov-GigaPath [xu2024whole] | ||||||||
| H-optimus-0 [hoptimus0] | ||||||||
| StainNet-Small (Ours) | ||||||||
| StainNet-Base (Ours) | ||||||||
@@
| MIST | Acc. | Bal acc. | AUC | F1-score | Acc. | Bal acc. | AUC | F1-score |
| ResNet-50 [he2016deep] | ||||||||
| CTransPath [wang2022transformer] | ||||||||
| PathoDuet [hua2024pathoduet] | ||||||||
| HIPT [chen2022scaling] | ||||||||
| Lunit [kang2023benchmarking] | ||||||||
| UNI [chen2024towards] | ||||||||
| PathOrchestra [yan2025pathorchestra] | ||||||||
| GPFM [ma2025generalizable] | ||||||||
| UNI-2 [chen2024towards] | ||||||||
| Virchow-2 [zimmermann2024virchow2] | ||||||||
| Prov-GigaPath [xu2024whole] | ||||||||
| H-optimus-0 [hoptimus0] | ||||||||
| StainNet-Small (Ours) | ||||||||
| StainNet-Base (Ours) | ||||||||
4.3 Comparison results on ROI tasks
Table 3 and Table 4 present the comparative results on the IHC tasks and the special stain glomerulus tasks. For the IHC tasks, we observe that Virchow-2, which is pre-trained on both H&E and IHC images, and StainNet-Base both achieve superior overall performance compared to other models. However, considering that Virchow-2 is based on a ViT-Huge/14 architecture, its computational cost is substantially higher than that of StainNet-Base. Meanwhile, StainNet-Small also demonstrates a strong balance between computational efficiency and performance when compared with models of similar or even larger scales. It is worth noting that our goal is not to directly replace previous PFMs, but rather to highlight the advantages of pre-training on domain-specific pathology images. We further believe that training on larger-scale non-H&E datasets and using models with bigger parameter size will lead to further performance gains, which we leave as an important direction for future work. For the special stain tasks, we find that the advantage of StainNet-Small diminishes when compared with models of similar scale; however, StainNet-Base further widens the performance gap with other larger PFMs. We also provide Umap visualization for BCI in Appendix D.
| Patch encoder | Glomerulus-Masson | Glomerulus-PAS | Glomerulus-PASM | |||
| Bal acc. | F1-score | Bal acc. | F1-score | Bal acc. | F1-score | |
| ResNet-50 [he2016deep] | ||||||
| CTransPath [wang2022transformer] | ||||||
| PathoDuet [hua2024pathoduet] | ||||||
| HIPT [chen2022scaling] | ||||||
| Lunit [kang2023benchmarking] | ||||||
| UNI [chen2024towards] | ||||||
| UNI-2 [chen2024towards] | ||||||
| Virchow-2 [zimmermann2024virchow2] | ||||||
| Prov-GigaPath [xu2024whole] | ||||||
| H-optimus-0 [hoptimus0] | ||||||
| StainNet-Small (Ours) | ||||||
| StainNet-Base (Ours) | ||||||
@@ 4.4 Comparison of image retrieval
We also examine the retrieval capability of StainNet-Small and StainNet-Base to assess their semantic representation quality on special staining histology images. We compare StainNet with CTransPath and UNI on the MIST dataset, as shown in Figure 2. We observe that both StainNet models consistently achieve the best Recall@K and mAP@K across all choices of K, indicating that they have learned a well-structured embedding space for special staining images.
@@ 4.5 Effectiveness of different fine-tuning data ratios
We further evaluate the fine-tuning performance of StainNet models under varying amounts of downstream special-stain training data to assess their learning capability. Figure 3 presents the results of StainNet-Small, StainNet-Base, CTransPath, and UNI on the MIST dataset using linear and MLP probes under different proportions of fine-tuning samples. We observe that StainNet-Base achieves the best performance across all ratio settings. Although StainNet-Small performs slightly worse than UNI under the 10% data setting with a linear probe, it achieves the best performance in all other settings and substantially outperforms the CTransPath of similar parameter size. Notably, applying an MLP probe significantly increases the performance gap between StainNet and UNI in low-fine-tuning data scenarios.
4.6 Comparison results on H&E-stained tasks
We further explore the potential of the proposed StainNet models on H&E image tasks. Table 5 presents the quantitative results. We find that StainNet-Small is not significantly different from other PFMs with similar parameter sizes on both tasks, and that StainNet-Base does not improve performance on H&E image tasks despite having a larger parameter size than StainNet-Small. However, the large PFM UNI, specifically trained on H&E-stained images, leads in both task types, especially on the WSI task, where the gap is significant.
| Datasets | CRC-100K | KatherMS | PANDA-Karo. |
| HIPT [chen2022scaling] | |||
| CTransPath [wang2022transformer] | |||
| Lunit [kang2023benchmarking] | |||
| UNI [chen2024towards] | |||
| StainNet-Small (Ours) | |||
| StainNet-Base (Ours) |
4.7 Comparison of different pre-training data ratios and iterations
We also conduct an ablation study to investigate the impact of different pre-training data scales. Specifically, we randomly sample 25 and 50 patches from each WSI to construct two pre-training datasets of different sizes, which are then used to pretrain StainNet-Small with DINO. Then we adapt the pre-trained models to the BCI and MIST ROI-level downstream tasks, as shown in Figure 4.a. We observe that larger pre-training datasets lead to stronger downstream performance, which validates the scalability of DINO on non-H&E images.
Finally, we investigate the performance variation of StainNet-Small under different pre-training iterations and the behavior of its loss curve. We believe these observations can provide valuable insights for the development of SSL methods in computational pathology, where pre-training details such as the expected loss curve and convergence behavior are often underreported. Figure 4.b and Figure 4.c illustrate the pre-training loss curve over 100 epochs and the corresponding linear probe performance on the MIST and Glomerulus-PAS dataset at six iteration checkpoints. We observe that on MIST, the loss plateaus between epoch 30 and 50, but with significant performance improvement during this period. After epoch 60, the overall loss begins to decrease again, but with slight changes in performance. Ultimately, the model reaches its optimal state at the final iteration. On Glomerulus-PAS, performance increases progressively with the number of epochs. These findings suggest that the shift in the DINO pre-training loss may serve as a valuable indicator of representation quality. While additional exploration is required to generalize these observations to other settings, we encourage future SSL development in computational pathology to consider allocating more training epochs, which may be crucial for achieving optimal model performance.
@@ 5 Conclusion and future works
In this work, we propose StainNet, a collection of pathology foundation models specifically for special staining histopathology image representation. Utilizing the DINO self-supervised framework, we train StainNet-Small and StainNet-Base models based on over 1.4 million unlabeled images from non-H&E WSIs in HISTAI. We evaluate StainNet models on three in-house IHC WSI tasks, two public IHC ROI tasks, and three in-house special stain ROI tasks, including classification, few-ratio settings, and image retrieval. The results demonstrate that StainNet models learn strong and robust representations for IHC and special stain pathology images.
For future work, we plan to develop larger StainNet models by expanding the pre-training dataset, scaling the model parameter size, and evaluating them on broader downstream tasks. We will also explore integrating StainNet with H&E-based PFMs for multimodal applications such as survival prognosis and therapy response prediction. We believe that StainNet will further accelerate the research and applications in computational pathology involving diverse histological modalities.
@@
Appendix A Details of pre-training datasets
We observe that the non-H&E WSI data collected in HISTAI exhibit a certain degree of ambiguity in naming conventions. Specifically, heterogeneous clone identifiers and inconsistent file naming rules are used across different cases, and some WSIs lack explicit annotations of their corresponding non-H&E staining types. As a result, it is practically challenging to assign fully accurate and unified stain labels to all WSIs. Nevertheless, we made systematic efforts to curate and organize the data to the best extent possible, with the goal of providing the community with a reasonably accurate distribution for the pre-training data, thereby reflecting the overall composition and diversity of non-H&E staining modalities. Based on our statistics, the HISTAI non-H&E pre-training dataset comprises 331 distinct immunohistochemistry (IHC) staining types and 12 special staining types. Figure 7 illustrates the overall distribution of different staining modalities within the pre-training data. Furthermore, Table 6 reports the distribution of IHC staining types with more than 20 cases, highlighting the major IHC sources contributing to the pre-training process. Table 7 summarizes the case distribution of all special staining types. We hope that these statistics and visualizations will serve as a useful reference for future studies leveraging HISTAI non-H&E data.
| IHC stain type | Case number | Patch image number |
| Ki67 | 1419 | 95715 |
| ER | 762 | 51533 |
| PR | 691 | 46757 |
| CD20 | 535 | 39513 |
| CD138 | 464 | 41006 |
| HER2/neu | 441 | 22850 |
| MCK | 431 | 28564 |
| CD3 (VET) | 403 | 29128 |
| CD10 | 385 | 35373 |
| CK7 | 365 | 25864 |
| Bcl-2 | 313 | 28075 |
| Bcl-6 | 311 | 27345 |
| S100 (4c4) | 289 | 24579 |
| CD23 | 287 | 24359 |
| CD34 | 282 | 23313 |
| PD-L1 | 281 | 25500 |
| Synaptophysin | 272 | 18537 |
| CD5 | 269 | 20015 |
| P63-7JUL | 253 | 14431 |
| HER2 | 249 | 8607 |
| Cyclin-D1 | 247 | 22386 |
| MUM1-Protein | 233 | 20339 |
| TTF-1 | 230 | 13932 |
| CK20 (ks20) | 214 | 14266 |
| CD30 | 208 | 17529 |
| SMA | 202 | 18070 |
| PAX-8 | 190 | 14440 |
| Chromogranin-A | 185 | 13375 |
| CD45 | 183 | 16063 |
| GATA-3 | 178 | 11796 |
| CDX2 | 174 | 10253 |
| Desmin | 154 | 10415 |
| MSH6 | 154 | 10044 |
| MlH1 | 151 | 11546 |
| SOX-10 | 142 | 14610 |
| MSH2 | 141 | 11224 |
| PAX-5 | 138 | 10074 |
| PMS2 | 136 | 10648 |
| CD56 | 128 | 9803 |
| Vimentin | 126 | 11039 |
| Melan-A | 124 | 10128 |
| IHC stain type | Case number | Patch image number |
| P53 | 123 | 9530 |
| P16 | 119 | 7881 |
| HMB45 | 116 | 9334 |
| CD56 (123c3) | 112 | 7871 |
| P40 | 112 | 6112 |
| AMACR | 112 | 4823 |
| Epstein-Barr-Birus | 111 | 10141 |
| CK14 | 104 | 5918 |
| CD117 | 100 | 8255 |
| WT1-WT49 | 100 | 7267 |
| CD68 | 99 | 8079 |
| E-Cadherin | 92 | 4808 |
| CK5&6 | 90 | 5444 |
| S100 | 81 | 7301 |
| CD15 | 79 | 6646 |
| CD4 | 70 | 5575 |
| Kappa-Light-Chain | 67 | 5637 |
| Lambda-Light-Chain | 65 | 4974 |
| EMA (gp1) | 63 | 5318 |
| CD8 | 61 | 4992 |
| A1K | 59 | 4542 |
| DOG-1 | 58 | 4482 |
| CD79A | 56 | 4088 |
| CD99 | 54 | 5074 |
| Calponin-1 | 54 | 4412 |
| Calretinin | 52 | 4395 |
| Arginase-1 | 49 | 4299 |
| SATB2 | 49 | 3918 |
| CD16 | 47 | 4400 |
| ERG | 47 | 3070 |
| HlA-DR-Antigen | 46 | 3525 |
| Glypican-3 | 46 | 3104 |
| CK8 & CK18 (b22) | 46 | 2957 |
| Napsin-a | 46 | 2906 |
| INSML | 43 | 3268 |
| CK-HMW | 43 | 1532 |
| GFAP | 42 | 3566 |
| P63 | 42 | 3167 |
| CD31 | 40 | 3207 |
| CD2 | 38 | 3012 |
| Beta-Catenin | 37 | 3231 |
| ER-other | 37 | 3037 |
| CD21 | 36 | 2869 |
| IHC stain type | Case number | Patch image number |
| INI-1 | 35 | 3383 |
| CD7 | 35 | 2468 |
| TDT | 32 | 2934 |
| CK19 | 32 | 2278 |
| Androgen-Receptor | 32 | 1747 |
| Inhibin-Alpha | 29 | 2646 |
| Mammaglobin | 29 | 2089 |
| PSA | 29 | 1796 |
| C-MYC | 27 | 2393 |
| HHV8 | 27 | 2067 |
| CD3-K | 27 | 1636 |
| Myeloperoxidase | 26 | 2424 |
| CD61 | 26 | 2096 |
| CK20 (KS20) | 26 | 2071 |
| STAT6 | 26 | 1992 |
| MDM2 | 26 | 1850 |
| CD38 | 26 | 1643 |
| Heppar | 25 | 2213 |
| PLAP | 25 | 1968 |
| LIFR | 23 | 2249 |
| IGG4 | 23 | 2106 |
| Caix-TH22 | 23 | 2059 |
| CK-5-6-d5-16b4 | 23 | 1653 |
| CEA | 22 | 1677 |
| Alpha-Fetoprotein | 21 | 1850 |
| SAll4 | 21 | 1504 |
| CD3 | 21 | 1453 |
| Podoplanin | 20 | 1845 |
| Total | 16150 | 1216942 |
| Special stain type | Case number | Patch image number |
| Giemsa | 869 | 114016 |
| Argentum | 67 | 2956 |
| Van Gieson | 38 | 5039 |
| PAS | 30 | 3061 |
| Alcian Blue | 31 | 3743 |
| Congo Red | 17 | 763 |
| Toluidine Blue | 16 | 1011 |
| Calcitonin | 9 | 816 |
| Ziehl Neelsen | 8 | 881 |
| PAS-D | 1 | 100 |
| Romanovsky | 1 | 100 |
| Grocott | 1 | 14 |
| Total | 1088 | 132500 |
Appendix B Details of DINO pre-training settings
We provide the data augmentation strategy and detailed hyperparameter settings for DINO, as shown in Table 8 and Table 9.
| Augmentation | Value |
| Image size | |
| Local view size | |
| Global random resize scale | (0.4, 1.0) |
| Local random resize scale | (0.05, 0.4) |
| Number of local views | 8 |
| Random horizontal flip | 0.5 |
| Random vertical flip | 0.0 |
| Random gray scale | 0.2 |
| Color jitter brightness | 0.8 |
| Color jitter contrast | 0.8 |
| Color jitter hue | 0.2 |
| Color jitter prob | 0.8 |
| Color jitter saturation | 0.4 |
| Color jitter strength | 0.5 |
| Gaussian blur limit | 0 |
| Gaussian blur prob | 1.0 |
| Gaussian blur sigmas | (0.1, 2.0) |
| Global view Gaussian blur prob | 0.1 |
| Global view Solarize prob | 0.2 |
| Global view Solarize threshold | 0.5 |
| Local view: Gaussian blur prob | 0.5 |
| Normalization mean | (0.5, 0.5, 0.5) |
| Normalization std | (0.5, 0.5, 0.5) |
We provided some examples of original images and their augmented versions, which are illustrated in Figure 6.
| Hyperparameter | Value |
| Batch normalization | False |
| Bottleneck dimension | 256 |
| Center momentum | 0.9 |
| Hidden dimension | 2048 |
| Learning rate scale method | linear |
| Momentum end | 1.0 |
| Momentum start | 0.9995 |
| Last layer normalization | True |
| Output dimension | 65536 |
| Epoch | 100 |
| Total batch size | 256 |
| Student freeze last layer epochs | 0 |
| Student temperature | 0.1 |
| Teacher temperature | 0.04 |
| Warmup teacher temperature | 0.04 |
| Warmup teacher temperature epochs | 0 |
| Weight decay end | 0.4 |
| Weight decay start | 0.04 |
| Optimizer | AdamW |
| Optimizer betas | (0.9, 0.999) |
| Optimizer eps | 1e-8 |
| Optimizer learning rate | 5e-4 |
| Optimizer weight decay | 0.04 |
| Automaticmixedprecision | BF16 |
Appendix C Statistical significance analysis
| Model | Acc. () | Bal acc. () | AUC () | F1-score () |
| ResNet-50 | 2.70e-04*** | 3.34e-03** | 1.16e-03** | 1.03e-04*** |
| CTransPath | 2.40e-01 | 3.13e-02* | 1.36e-01 | 3.85e-02* |
| PathoDuet | 8.26e-03** | 2.57e-02* | 2.37e-03** | 5.93e-03** |
| HIPT | 4.66e-02* | 2.48e-02* | 3.08e-02* | 4.43e-02* |
| Lunit | 3.27e-02* | 2.85e-02* | 8.04e-04*** | 7.86e-03** |
| UNI | 9.94e-01 | 6.03e-02 | 4.93e-01 | 1.40e-01 |
| PathOrchestra | 3.68e-02* | 5.65e-02 | 2.23e-02* | 3.61e-02* |
| GPFM | 9.94e-01 | 6.03e-02 | 1.77e-01 | 1.44e-01 |
| UNI-2 | 8.96e-02 | 3.40e-02* | 7.33e-03** | 6.78e-03** |
| Virchow-2 | 3.33e-02* | 3.89e-02* | 6.68e-02 | 3.13e-03** |
| Prov-GigaPath | 1.61e-02* | 1.12e-01 | 4.99e-01 | 6.55e-02 |
| H-optimus-0 | 1.78e-01 | 4.66e-02* | 9.12e-02 | 4.50e-02* |
| StainNet-Small | 9.83e-02 | 2.60e-02* | 2.10e-03** | 1.72e-02* |
Appendix D Feature visualization